Статьи в реферируемых журналах:

1. M.A.Berezovskii, A.I.Dyachenko, I.K.Konkashbaev, V.B.Lopatko, Yu.V.Medvedev, I.V.Petrov and A.M.Rubenchik, Interaction of counterstreaming plasmas, Plasma Physics and Controlled Fusion,пар 26, # 12B, pp. 1477-1488 (1984).

2. М.А. Березовский, А.И. Дьяченко, А.М. Рубенчик, О взаимодействии электронного пучка с сильнотурбулентной плазмой, ЖЭТФ, 88, Вып. 4, стр. 1191-1196 (1985).

3. А.И. Дьяченко, В.Е. Захаров, А.М. Рубенчик, Р.З Сагдеев, В.Ф. Швец, Двумерный ленгмюровский коллапс и двумерные ленгмюровские кавитоны, Письма в ЖЭТФ, 44, Вып. 11, стр. 504-507 (1986).

4. M.A.Berezovskii, A.I.Dyachenko, I.V.Petrov, A.N.Pushkarev, and A.M.Rubenchik A.M., On Powerful Electron Beam Relaxation in Magnetized Plasma, Phys. Lett. A, 121, pp. 178-181 (1987).

5. А.И. Дьяченко, В.Е. Захаров, А.М. Рубенчик, Р.З Сагдеев, В.Ф. Швец, Численное Моделирование двумерного ленгмюровского коллапса, ЖЭТФ, 94 Вып. 3, стр. 144-155 (1988).

6. А.И. Дьяченко, В.Е. Захаров, А.Н. Пушкарев, В.Ф. Швец, В.В. Яньков, Солитонная турбулентность в неинтегрируемых волновых системах, ЖЭТФ, 96, Вып. 6(12), стр. 2026-2031, (1989).

7. А.И. Дьяченко, А.Н. Пушкарев, А.М. Рубенчик, Численное моделирование взаимодействия электронных пучков с плазмой, Физика плазмы, 15, стр. 790-796 (1989).

8. A.I.Dyachenko, A.N.Puchkarev, A.M.Rubenchik, V.F.Shvets, A Particle Model for Three-Dimensional Langmuir Collapse Simulation, Computer Physics Communications, 60, pp. 239-245 (1990).

9. A.I.Dyachenko, A.N.Pushkarev, A.M.Rubenchik, R.Z.Sagdeev, V.F.Shvets, and V.E.Zakharov, Computer Simulation of Langmuir Collapse, Physica D, 52, pp. 78-102 (1991).

10. A.I.Dyachenko, A.C.Newell, A.N.Pushkarev, and V.E.Zakharov, Optical Turbulence, Physica D, 57, pp. 96-160 (1992)

11. A.I.Dyachenko, S.V.Nazarenko, and V.E.Zakharov, Wave-Vortex Dynamics in the Drift and β-Plane Turbulence, Phys. Lett. A, 165, pp. 330-334 (1992)

12. A.I.Dyachenko, E.A.Kuznetsov, Instability and Self-Focusing of Solitons in the Boundary Layer, Письма в ЖЭТФ, 59, Вып. 2, стр. 103-108 (1994).

13. A.I.Dyachenko, V.E.Zakharov, Is the free-surface hydrodynamics an integrable system? Phys. Lett. A 190, pp. 144-148 (1994).

14. A.I.Dyachenko, Y.V.L’vov, On the Hasselmann’s and Zakharov’s approaches to the kinetic equations for the gravity waves J.Phys. Oceanography 25 N12, p. 3237-3238, (1995)

15. A.I.Dyachenko, Y.V.L’vov and V.E.Zakharov, Five-wave interaction on the surface of deep fluid, Physica D 87, pp. 233-261, (1995)

16. A.I.Dyachenko, E.A.Kuznetsov, Two-Dimentional Wave Collapse in The Boundary Layer, Physica D 87, pp. 301-313, (1995)

17. Finn V.S., Baker V.R., Dolginov A.Z., Gabitov I. and Dyachenko A.I., Large scale spatial patterns in topography at Alpha Regio, Venus. Geophysical Research Letters, 22 no.14, pp. 1901-1904, (1995)

18. A.I.Dyachenko, E.A.Kuznetsov, M.D.Spector and V.E.Zakharov, Analytical Description of the Free Surface Dynamics of an Ideal Fluid. Phys. Lett. A, 221, pp. 73-79, (1996).

19. A.I.Dyachenko, V.E.Zakharov, Toward an Integrable Model of Deep Water. Phys. Lett. A, 221 (1996) 80-87.

20. А.И. Дьяченко, В.Е. Захаров, Е.А. Кузнецов, Нелинейная динамика свободной поверхности идеальной жидкости, Физика плазмы, 22, Вып. 10, стр. 916-928 (1996)

21. A.I.Dyachenko, G.Fal’kovich, Condensate Turbulence in Two Dimensions. Physical Review E, 54, # 5, pp. 5095-5099, (1996)

22. V.E.Zakharov, A.I.Dyachenko, High-Jacobian approximation in the free surface dynamics of an ideal fluid. Physica D 98, pp. 652-664, (1996).

23. А.И. Дьяченко, О динамике идеальной жидкости со свободной поверхностью. Доклады Академии Наук, 376, Вып. 1, стр. 27-29, (2001).

24. V.E. Zakharov, O.A. Vasilyev and A.I. Dyachenko, Kolmogorov spectra in one-dimensional weak turbulence, Письма в ЖЭТФ, 73, Вып. 2, стр. 68-70 25. V.E. Zakharov, A.I. Dyachenko and O.A. Vasilyev, New method for numerical simulation of a nonstationary potential flow of incompressible fluid with a free surface. European Journal of Mechanics B-Fluids, 21, No. 3, pp. 283-291, (2002).

26. M. Chertkov, Y.-J. Chung, A. Dyachenko, I. Gabitov, I. Kolokolov, and V. Lebedev, Shedding and interaction of solitons in weakly disordered optical fibers. Phys. Rev. E, 67, No. 3, 036615, (2003).

27. A.I. Dyachenko, A.O. Korotkevich, V.E. Zakharov, Decay of the monochromatic capillary wave. Письма в ЖЭТФ, 77, Вып. 9-10, стр. 572-576 (2003).

28. A.I. Dyachenko, A.O. Korotkevich, V.E. Zakharov, Weak Turbulence of Gravity Waves. Письма в ЖЭТФ, 77, Вып. 9-10, стр. 649-653 (2003).

29. A.I. Dyachenko, A.O. Korotkevich, V.E. Zakharov, Weak Turbulent Kolmogorov Spectrum for Surface Gravity Waves. Phys. Rev. Lett., 92, No. 13, pp. 134501-1–134501-4, (2004).